# A Comprehensive Study on Large Language Models for Mutation Testing

### v4

@@ 2.4.2. RQ2: Validity
We consider three key metrics to evaluate the validity of different mutation generation approaches: generation efficiency, compilability, and the proportion of duplicate and equivalent mutations, providing a basic understanding of the validity of each approach.
RQ2.1: Compilability Rate: the proportion of mutations that successfully compile, which is computed as:
@@
RQ2.2: Duplicate Mutation Rate: the proportion of duplicate mutations, computed as:
LLMs sometimes generate duplicate mutations, i.e., mutations that are either identical to the original code or identical to each other. While traditional approaches rarely generate duplicate mutations, LLMs can produce them due to their probabilistic text-generation mechanisms and learned patterns from large-scale training data. After eliminating duplicate mutations, all remaining mutations are guaranteed to be syntactically distinct both from the original code and from each other.
RQ2.3: Equivalent Mutation Rate: the proportion of equivalent mutation, computed as:
Equivalent mutations are semantically equivalent to the original program, which impacts the accuracy of the mutation score. We prefer the approach with a lower EMR. Note that while equivalent mutations are also considered semantically duplicate, they are not easily identifiable. Therefore, we distinguish them from other duplicate mutations, which can be filtered syntactically.
@@ 2.4.3. RQ3: Efficiency
Efficiency describes how fast a mutation approach produces mutations. Mutation testing requires numerous mutations to evaluate the robustness of a test suite effectively. Therefore, the efficiency of mutation generation directly impacts the feasibility and scalability of mutation testing. To assess this, we compare these approaches based on time costs.
RQ3.1: Average Generation Time: measures the average time required to generate each mutation, providing a quantitative evaluation of the efficiency of each approach.
RQ3.2: Average Token: measures the average number of consumed tokens per mutation, reflecting both the financial cost (due to token-based pricing) and the computational time cost (since longer token sequences typically require more processing time). Since input and output token costs differ across models, we calculate them separately.
2.5. Mutation Generation Setups
In this subsection, we discuss the setups for mutation generation approaches, especially the setups for LLM-based generation.
For the well-established mutation generation approaches (i.e., PIT, Major, LEAM, and BERT), we activate all their mutation operators. Since we try to compare the similarity between the generated mutations and real bugs, we filter the mutations within the context of real-world bugs. Note that we supply the same code snippets as context across all mutation generation approaches to ensure a fair comparison.
@@ 3. Evaluation Results
In this section, we aim to comprehensively evaluate these approaches via various aspects. Table 4 presents the overall performance of each generation approach. In terms of the number of mutations generated (Mutation Count), PIT produces the largest number of mutations at 340,728, significantly more than other approaches, while the remaining approaches range from about twenty thousand to thirty thousand. In terms of mutation score, the traditional rule-based approaches (PIT and Major) are over 0.5, LEAM and BERT are over 0.6, and the LLMs from the GPT family are all over 0.7.
Note that the number of generated mutations and mutation scores are not quality metrics for evaluating these approaches. A large number of mutations may include redundant or equivalent mutants that do not contribute to mutation scores, while a high mutation score may be influenced by the ease with which mutants are detected rather than their resemblance to real faults.
@@
GPT-4o-M: GPT-4o-Mini. DS-761b: DeepSeek-V3-761b. DS-236b: DeepSeek-Coder-V2-236b. SC-16b: StarChat--16b. CL-13b: CodeLlama-Instruct-13b.
@@ 3.1. RQ1: Performance on Effectiveness
To measure the effectiveness of mutation generation approaches, we evaluate them with three sub-RQs, which estimate their effectiveness via real bug detectability, coupling, and diversity, respectively.
@@
Real Bug Detection Rate: The column Real Bug Detec. of Table 4 presents the proportion of real bugs that can be detected across different mutation generation approaches.
@@
For LLM-based approaches, the results are significantly more effective than those of traditional ones. The existing LLM-based approaches, BugFarm and LLMorpheus, achieve a rate of 58.6% and 68.5%, respectively. For the LLMs based on our prompting, DeepSeek-V3-761b achieves the highest detection rate at 91.1%, closely followed by GPT-4o-Mini, GPT-4o and DeepSeek-Coder-V2-236b, with detection rates of 90.8%, 89.9%, and 89.4%, respectively. Meanwhile, GPT-3.5 and CodeLlama achieve moderate detection rates of 87.0% and 71.2%, respectively.
| Proj. | Chart | Lang | Time | Math | Moc. | Clo. | Cli | Co. | Gs. | Csv | Ja. | Js. | ConD. | All |
@@
To better understand the detection of the generation approaches, we analyze their project-wise detectability, shown as Table 5. For individual projects, LLM-based approaches consistently achieve top performance across multiple projects. Notably, DeepSeek-V3-761b detects the largest number of bugs in most Defects4J projects and achieves the highest overall bug detection count. Similarly, DeepSeek-Coder-V2-236b, GPT-4o, and GPT-4o-Mini demonstrate strong performance, achieving the best results in projects such as Mockio, Csv, and ConDefects.
@@
Average Ochiai Coefficient: The column Avg. Ochiai of Table 4 presents the Ochiai Coefficient values across the approaches, which measures the semantic distance between mutations and real bugs. Among the traditional approaches, Major achieves an Ochiai coefficient of 35.5%, higher than the others. LEAM and BERT achieve coefficients of 32.4% and 29.3%, respectively, indicating that the best-performing rule-based approach, Major, outperforms both of them. Existing other LLM-based approaches achieve an overall lower Ochiai coefficient of 17.7% and 24.1%, respectively. The LLM-based approaches using our prompt achieve notably higher Ochiai coefficients. DeepSeek-V3-761b attains the highest value at 59.8%, followed by GPT-4o at 56.2% and GPT-4o-Mini at 54.1%. The LLMs based on our prompt outperform other existing methods, not only traditional approaches but also existing LLM-based ones, showing their superiority in generating mutations that are semantically closer to real bugs.
High Ochiai Bug Count: Following the study of Ojdanic et al. (Ojdanic et al., 2023), the mutation set with an Ochiai coefficient over 0.8 is considered to be of high semantic similarity with its corresponding real bug. For each project, we count the number of bugs that achieved a high Ochiai coefficient by each approach, shown as Table 6. The overall trend is similar to Real Bug Detectability, which indicates that the LLM DeepSeek-V3-761b, based on our prompt, achieves the highest number of bugs across most projects. The closed-source LLMs GPT-4o and GPT-4o-Mini also exhibit strong performance, closely following DeepSeek-V3-761b in many cases. Compared with our approach, all traditional approaches exhibit significantly fewer high Ochiai bugs. Interestingly, among traditional approaches, Major outperforms LEAM and BERT across most projects.
| Proj. | Chart | Lang | Time | Math | Moc. | Clo. | Cli | Co. | Gs. | Csv | Ja. | Js. | ConD. | All |
@@
Coupling Rate: The overall Coupling Rate for each mutation generation approach is shown in the last row of Table 4. The traditional approaches, PIT, Major, LEAM, and BERT, exhibit lower coupling rates than LLM-based approaches, with coupling rates of 23.5%, 38.9%, 38.1%, and 47.3%, respectively. Among the LLM-based approaches, LLMorpheus achieves the highest coupling rate, with a value of 51.7%. It is followed closely by DeepSeek-V3-761b and GPT-4o-Mini based on our prompt, with coupling rates of 50.8% and 50.6%, respectively. Overall, the trends are similar to the behavior similarity discussed in Section 3.1, indicating the superiority of the LLM-based approaches, particularly LLMorpheus, DeepSeek-V3-761b, and GPT-4o, in generating overall highly coupled mutations that better mimic real bugs.
@@
Coupling Categorization: As mentioned in Section 2.4.1, we classify each mutation of each generation approach into different scales, i.e., Strong Substitution, Strong + Additional Tests Fail, Partial Substitution, Partial + Additional Tests Fail, Not Substitution, and Not Detected. Note that the first four categories are used to calculate the Coupling Rate, indicating these mutations semantically overlap with their corresponding real bugs. Table 7 presents the distribution of each generation approach. Among the traditional approaches, PIT and Major exhibit a significantly higher proportion of mutants in the Not Detected category, with rates of 45.9% and 47.9%, respectively, indicating that a substantial number of their generated mutants fail to couple with real bugs. However, they achieve relatively higher proportions in the Strong category, with PIT at 10.1% and Major at 9.6%, suggesting that while their coupling ability is limited, they can still generate a small number of strongly coupled mutations. In contrast, LLM-based approaches such as DeepSeek-V3-761b, GPT-4o, and GPT-4o-Mini, demonstrate stronger coupling capabilities across multiple categories. For instance, DeepSeek-V3-761b achieves the highest proportion in the Strong category at 11.2%, while GPT-4o-Mini leads in the Strong + Additional Tests Fail category, with 12.1%. Overall, while the traditional rule-based approaches generate a larger proportion of strong substitution mutations, LLM-based approaches demonstrate a stronger ability to generate diverse and highly coupled mutations.
@@
We explore the syntactic diversity of mutations generated by each approach. Note that PIT works at the Java byte code level, we thus skip it in this RQ.
AST Node Diversity: To assess the diversity of AST node types, we first determine if a mutation involves deletion (i.e., removing a code element). For the non-deletion mutations, we parse the code before and after mutation by the parser Javalang555https://pypi.org/project/javalang, to observe which new AST node types are introduced. A larger number of newly introduced AST node types indicates richer diversity.
Figure 2 illustrates the number of introduced AST node types, with traditional approaches highlighted in green and LLM-based approaches in blue. The figure reveals that LLM-based mutation approaches introduce more types of new AST nodes than traditional approaches. The best LLM-based approaches, DeepSeek-V3-761b and DeepSeek-Coder-V2-236b, exhibit the greatest diversity, introducing 49 new AST node types, followed by GPT-4o-Mini, which introduces 45 new types. Additionally, all six LLMs based on our prompt are significantly higher than all other approaches in AST node types. The existing LLM-based approaches, BugFarm and LLMorpheus, introduce 31 and 21 new AST node types, respectively. While BugFarm outperforms traditional approaches, it still lags behind ours. We also check the AST node diversity of the examples we provided in the LLM prompts (see Table 2), as shown by the “Examples” in Figure 2. Interestingly, we find that these examples introduce only one new type of AST node and are subsumed by all other approaches. This indicates that LLMs are creative in generating mutations, producing significantly more diverse mutations than the examples provided to them.
@@
We further analyzed the distribution of deletions and AST node types across different approaches, as shown in Table 8, which highlights the proportion of the mutations deleting the target code element and the top three most frequent AST node types. For the proportion of deletions, the traditional approach, LEAM, shows the highest deletion mutation ratio, with 48.9%, followed by Major at 15.3%. In contrast, all LLM-based approaches generate deletion mutations at a rate below 6%, with three models producing less than 1.0%. This indicates that, compared to existing approaches, LLMs are less inclined to generate deletion mutations. As shown later, we find that deletions are one of the key factors contributing to the lower compilation success rate.
Regarding AST node types, the most frequent new AST node type of all the LLM-based approaches is Literal (e.g., ). However, for the traditional approach Major, the most frequent node type is Statement (e.g., replacing a statement with an empty statement). Note that although these mutations are not with the deletion operator, a significant portion of the production of empty statements exists, leading to a large proportion.
AST Edit Distance: To quantify the extent of syntactic modifications introduced by each mutation generation approach, we compute the AST edit distance for each mutation using Gumtree (Falleri et al., 2014). The distribution of edit distances for each approach is presented in Table 9, providing insights into the complexity and diversity of the generated mutations.
@@
From the table, we can find that all traditional approaches predominantly generate small-scale changes. Specifically, the proportion of mutations with an edit distance of 1 for Major, LEAM, and BERT is 69.2%, 58.5%, and 73.8%, respectively, indicating they involve minimal syntactic changes to the original code in the majority of cases. In contrast, LLM-based approaches exhibit a broader range of edit distances, reflecting their capability to introduce more complex transformations. For example, DeepSeek-V3-761b exhibits a more balanced distribution, with 44.2% of its mutations having an edit distance of 2, 7.2% with a distance of 3, and 2.2% with a distance of 4, indicating it generates both minor and moderate code changes.

### v5

@@ 2.4.2. RQ2: Validity
We consider three key metrics to evaluate the validity of different mutant generation approaches: generation efficiency, compilability, and the proportion of duplicate and equivalent mutants, providing a basic understanding of the validity of each approach.
RQ2.1: Compilability Rate: the proportion of mutants that successfully compile, which is computed as:
@@
RQ2.2: Duplicate Mutant Rate: the proportion of duplicate mutants, computed as:
LLMs sometimes generate duplicate mutants, i.e., mutants that are either identical to the original code or identical to each other. While traditional approaches rarely generate duplicate mutants, LLMs can produce them due to their probabilistic text-generation mechanisms and learned patterns from large-scale training data. After eliminating duplicate mutants, all remaining mutants are guaranteed to be syntactically distinct both from the original code and from each other.
RQ2.3: Equivalent Mutant Rate: the proportion of equivalent mutants, computed as:
Equivalent mutants are semantically equivalent to the original program, which impacts the accuracy of the mutation score. We prefer the approach with a lower EMR. Note that while equivalent mutants are also considered semantically duplicate, they are not easily identifiable. Therefore, we distinguish them from other duplicate mutants, which can be filtered syntactically.
@@ 2.4.3. RQ3: Efficiency
Efficiency describes how fast a mutation approach produces mutants. Mutation testing requires numerous mutants to evaluate the robustness of a test suite effectively. Therefore, the efficiency of mutant generation directly impacts the feasibility and scalability of mutation testing. To assess this, we compare these approaches based on time costs.
RQ3.1: Average Generation Time: measures the average time required to generate each mutant, providing a quantitative evaluation of the efficiency of each approach.
RQ3.2: Average Token: measures the average number of consumed tokens per mutant, reflecting both the financial cost (due to token-based pricing) and the computational time cost (since longer token sequences typically require more processing time). Since input and output token costs differ across models, we calculate them separately.
2.5. Mutant Generation Setups
In this subsection, we discuss the setups for mutant generation approaches, especially the setups for LLM-based generation.
For the well-established mutant generation approaches (i.e., PIT, Major, LEAM, and BERT), we activate all their mutation operators. Since we try to compare the similarity between the generated mutations and real bugs, we filter the mutants within the context of real-world bugs. Note that we supply the same code snippets as context across all mutant generation approaches to ensure a fair comparison.
@@ 3. Evaluation Results
In this section, we aim to comprehensively evaluate these approaches via various aspects. Table 4 presents the overall performance of each generation approach. In terms of the number of mutants generated (Mutant Count), PIT produces the largest number of mutants at 340,728, significantly more than other approaches, while the remaining approaches range from about twenty thousand to thirty thousand. In terms of mutation score, the traditional rule-based approaches (PIT and Major) are over 0.5, LEAM and BERT are over 0.6, while most LLM-based approaches are over 0.7. Particularly, BugFarm (DS-671b) achieves the highest mutation score, which is nearly 0.79.
Note that the number of generated mutants and mutation scores are not quality metrics for evaluating these approaches. A large number of mutants may include redundant or equivalent mutants that do not contribute to mutation scores, while a high mutation score may be influenced by the ease with which mutants are detected rather than their resemblance to real faults.
@@
GPT-4o-M: GPT-4o-Mini. DS-671b: DeepSeek-V3-671b. DS-236b: DeepSeek-Coder-V2-236b. SC-16b: StarChat--16b. CL-13b: CodeLlama-Instruct-13b.
@@ 3.1. RQ1: Performance on Effectiveness
To measure the effectiveness of mutant generation approaches, we evaluate them with three sub-RQs, which estimate their effectiveness via real bug detectability, coupling, and diversity, respectively.
@@
Real Bug Detection Rate: The column Real Bug Detec. of Table 4 presents the proportion of real bugs that can be detected across different mutant generation approaches.
@@
For LLM-based approaches, the results are significantly more effective than those of traditional ones. Across all models, LLMorpheus and LLMut constantly outperform all traditional approaches. For example, LLMut (DS-671b) achieves the highest detection rate at 91.1%, closely followed by LLMut (GPT-4o-M) and LLMut (GPT-4o), with detection rates of 90.8% and 89.9%, respectively. Moreover, we can find that all DS-671b outperforms GPT-4o in terms of bug detection rate, across all three LLM-based approaches.
| Proj. | Chart | Lang | Time | Math | Moc. | Clo. | Cli | Co. | Csv | Gs. | Ja. | Js. | ConD. | All |
@@
To better understand the detection of the generation approaches, we analyze their project-wise detectability, shown as Table 5. For individual projects, LLM-based approaches consistently achieve top performance across multiple projects. Notably, LLMut (DS-671b) detects the largest number of bugs in most Defects4J projects and achieves the highest overall bug detection count. Similarly, LLMut (DS-236b), LLMut (GPT-4o), and LLMut (GPT-4o-M) demonstrate strong performance, achieving the best results in projects such as Mockio, Csv, and ConDefects.
@@
Average Ochiai Coefficient: The column Avg. Ochiai of Table 4 presents the Ochiai Coefficient values across the approaches, which measure the semantic distance between mutants and real bugs. Among the traditional approaches, Major achieves an Ochiai coefficient of 35.5%, higher than the others. LEAM and BERT achieve coefficients of 32.4% and 29.3%, respectively, indicating that the best-performing rule-based approach, Major, outperforms both of them.
The average Ochiai coefficients of LLM-based approaches vary significantly, ranging from 17.7% to 59.8%, indicating diverse semantic distances of their mutants. Among them, LLMut (DS-671b) achieves the highest value at 59.8%, followed by LLMut (GPT-4o) at 56.2% and LLMut (GPT-4o-M) at 54.1%, indicating that these mutants capture fault-triggering behaviors more closely aligned with real bugs compared to traditional approaches.
High Ochiai Bug Count: Following the study of Ojdanic et al. (Ojdanic et al., 2023), the mutant set with an Ochiai coefficient over 0.8 is considered to be of high semantic similarity with its corresponding real bug. For each project, we count the number of bugs that achieved a high Ochiai coefficient using each approach, as shown in Table 6. The overall trend is similar to Real Bug Detectability, which indicates that the LLMut (DS-671b) achieves the highest number of bugs across most projects. Moreover, LLMut (GPT-4o) and LLMut (GPT-4o-M) also exhibit strong performance, closely following LLMut (DS-671b) in many cases. Compared with our approach, all traditional approaches exhibit significantly fewer high Ochiai bugs. Interestingly, among traditional approaches, the rule-based approach Major outperforms the other learning-based ones, LEAM and BERT, across most projects.
| Proj. | Chart | Lang | Time | Math | Moc. | Clo. | Cli | Co. | Csv | Gs. | Ja. | Js. | ConD. | All |
@@
Coupling Rate: The overall coupling rate for each mutant generation approach is shown in the column Coup. Rate of Table 4. The traditional approaches, PIT, Major, LEAM, and BERT, exhibit lower coupling rates than LLM-based approaches, with coupling rates of 23.5%, 38.9%, 38.1%, and 47.3%, respectively. Among the LLM-based approaches, LLMorpheus (DS-671b) and LLMorpheus (GPT-4o) achieve the highest coupling rates, reaching 52.0% and 51.7%, respectively, suggesting that LLMorpheus produces mutants whose failing behaviors align most closely with those of real bugs. LLMut (DS-671b) and LLMut (GPT-4o-M) closely follow LLMorpheus, with coupling rates of 50.8% and 50.6%, respectively. Overall, the trends are similar to the behavior similarity discussed in Section 3.1, indicating the superiority of the LLM-based approaches, particularly LLMorpheus (DS-671b), in generating overall highly coupled mutants that better mimic real bugs.
@@
Coupling Categorization: As mentioned in Section 2.4.1, we classify each mutant of each generation approach into different scales, i.e., Strong Substitution, Strong + Additional Tests Fail, Partial Substitution, Partial + Additional Tests Fail, Not Substitution, and Not Detected. Note that the first four categories are used to calculate the Coupling Rate, indicating these mutants semantically overlap with their corresponding real bugs.
Table 7 presents the distribution of each generation approach. Among the traditional approaches, PIT and Major exhibit a significantly higher proportion of mutants in the Not Detected category, with rates of 45.9% and 47.9%, respectively, indicating that a substantial number of their generated mutants fail to couple with real bugs. However, they achieve relatively higher proportions in the Strong category, with PIT at 10.1% and Major at 9.6%, suggesting that while their coupling ability is limited, they can still generate a small number of strongly coupled mutants. In contrast, LLM-based approaches overall demonstrate stronger coupling capabilities across multiple categories. For instance, LLMorpheus (DS-671b) achieves the highest proportion in the Strong category at 11.8%, closely followed by LLMut (DS-671b) with the value of 11.2%. BugFarm (DS-671b) also achieves a notable 10.0% of strongly coupled mutants. In terms of the Strong + Additional Tests Fail category, LLMut (GPT-4o-M) leads with a proportion of 12.1%. Additionally, BugFarm and LLMorpheus exhibit the lowest proportion of No Substitution and Not Detected, indicating their mutation strategies target to. Overall, LLM-based approaches generate a larger proportion of both strongly coupled and partially coupled mutants, demonstrating a stronger ability to generate diverse and highly coupled mutants.
@@
We explore the syntactic diversity of mutants generated by each approach. Note that PIT works at the Java bytecode level, we thus skip it in this RQ.
AST Node Diversity: To assess the diversity of AST node types, we first determine if a mutant involves deletion (i.e., removing a code element). For the non-deletion mutants, we parse the code before and after mutation by the parser Javalang555https://pypi.org/project/javalang, to observe which new AST node types are introduced. A larger number of newly introduced AST node types indicates richer diversity.
Figure 2 illustrates the number of introduced AST node types, with traditional approaches highlighted in green and LLM-based approaches in blue. The figure reveals that LLM-based approaches present more types of new AST nodes than traditional approaches. The best LLM-based approaches, LLMut (DS-671b) and LLMut(DS-236b), exhibit the greatest diversity, introducing 49 new AST node types, followed by GPT-4o-Mini, which introduces 45 new types. Additionally, all models with LLMut introduce significantly more AST node types than traditional approaches. Meanwhile, BugFarm (DS-671b) and BugFarm (GPT-4o) present 33 and 31 new AST node types, respectively, which also outperform other traditional approaches. We also check the AST node diversity of the examples we provided in the LLM prompts (see Table 2), as shown by the “Examples” in Figure 2. Interestingly, we find that these examples introduce only one new type of AST node and are subsumed by all other approaches. This indicates that LLMs are creative in generating mutants, producing significantly more diverse mutants than the examples provided to them.
@@
We further analyzed the distribution of deletions and AST node types across different approaches, as shown in Table 8, which highlights the proportion of the mutants that delete the target code element and the top three most frequent AST node types. For the proportion of deletions, the traditional approach, LEAM, shows the highest deletion mutant ratio, with 48.9%, followed by Major at 15.3%. In contrast, all LLM-based approaches generate deletion mutants at a rate below 6%, with three models producing less than 1.0%. This indicates that, compared to existing approaches, LLMs are less inclined to generate deletion mutants. As shown later, we find that deletions are one of the key factors contributing to the lower compilation success rate.
Regarding AST node types, the most frequent new AST node type of all the LLM-based approaches is Literal (LT) (e.g., ). However, for the traditional approach Major, the most frequent node type is Statement (e.g., replacing a statement with an empty statement). Note that although these mutants are not with the deletion operator, a significant portion of the production of empty statements exists, leading to a large proportion. For the second most frequent new node types, BinaryOperation (BO) appears most often among LLM-based approaches, indicating that LLMs frequently modify or insert binary expressions such as arithmetic or comparison operators. For the third most frequent category, LLM-based approaches introduce a variety of AST nodes, including MethodInvocation (MI), MemberReference (MR), and ClassCreator (CC), reflecting their broader syntactic diversity in mutant generation.
AST Edit Distance: To quantify the extent of syntactic modifications introduced by each mutant generation approach, we compute the AST edit distance for each mutant using Gumtree (Falleri et al., 2014). The distribution of edit distances for each approach is presented in Table 9, providing insights into the complexity and diversity of the generated mutants.
@@
From the table, we can find that all traditional approaches predominantly generate small-scale changes. Specifically, the proportion of mutants with an edit distance of 1 for Major, LEAM, and BERT is 69.2%, 58.5%, and 73.8%, respectively, indicating they involve minimal syntactic changes to the original code in the majority of cases. In contrast, LLM-based approaches exhibit a broader range of edit distances, reflecting their capability to introduce more complex transformations. For example, only 38.3% of the mutants of BugFarm (DS-671b) have an edit distance of 1, and LLMut (DS-671b) generates 45.1% of mutants with an edit distance of 1 and 44.2% with a distance of 2, with an additional 6.9% corresponding to distances of 3. This pattern suggests that LLM-based methods, unlike traditional approaches, are capable of performing multi-node or compositional edits, leading to a richer diversity of syntactic structures.